# Electrocardiogram Instruction Tuning for Report Generation

### 07 Mar 2024 v1

@@ 1 Introduction
Diagnosing heart diseases is of paramount importance, requiring the analysis of diverse multimodal data sources, which includes electrocardiogram (ECG) recordings and clinical reports. Cardiologists interpret these ECG recordings to generate comprehensive ECG reports for heart disease diagnosis, which is a complex and time-consuming process. Recently, artificial intelligence (AI) models have been used to facilitate ECG data analysis by categorizing ECG data (Chen et al., 2022a; Hu et al., 2023; Liu et al., 2023a; Li et al., 2023b). Despite these early efforts, the automatic generation of reports from ECG recordings remains an underexplored area. Unlike other AI-empowered medical report generation applications (e.g., radiology report), the primary challenges for ECG report generation stem from the distinct nature of ECG content. ECG reports, often comprising brief phrases that summarize signal patterns, contrast with detailed anatomical descriptions in radiology reports. This difference in content and semantic interpretation between imaging and signal data complicates the direct application of radiology-focused methods to ECG reports. Furthermore, there is still a lack of comprehensive benchmarks for evaluating the performance of ECG report generation.
To tackle these challenges, we introduce MEIT, a Multimodal ECG Instruction Tuning framework that extends the capabilities of LLMs in the cardiology context to generate ECG reports using ECG signals based on human instructions. Inspired by the versatility of LLMs (Achiam et al., 2023; Touvron et al., 2023b; Wan et al., 2023; Wang et al., 2024) in handling diverse language tasks simultaneously, we develop a specialized instruction tuning process for ECG report generation. MEIT aligns human instructions with ECG signals, enabling LLMs to generate professional-grade reports and exhibit zero-shot report generation capabilities under domain transfer scenarios across various continents and data collection devices. Specifically, leveraging publicly available ECG datasets, we construct a multimodal instruction dataset including ECG records, human instructions, and paired reports. Then, we propose an effective and efficient attention-based fusion method to integrate ECG and text representations in the latent space, enabling LLMs to understand signals for report generation without introducing additional training parameters in the attention layer.
@@ 3.2 ECG Report Generation Model
where $i$ is the current layer index. Different from gated-attention fusion in Flamingo (Alayrac et al., 2022) and Memorizing Transformer (Wu et al., 2022), or Q-former in BLIP-2 (Li et al., 2023a) that require the additional trainable parameters or design complex modules, we propose a simple yet effective concatenated-fusion method that directly injects the ECG embeddings with language context in the causal-attention (C-Attn) of the models. In our approach, each attention layer combines $\mathbf{H}_{e}$, generated from the ECG encoder and projector as a prefix condition, with $\mathbf{H}_{t}^{i}$, derived from the preceding layer. The fusion process is as follows:
@@ 6 Related Work
Medical Report Generation. Medical Report Generation can be categorized into three main approaches: (1) Template Selection and Generation, where methods like HRGR (Li et al., 2018) focus on sentence creation or template retrieval, and CMAS (Jing et al., 2017) uses dual writers for differentiating regions; (2) Data Integration and Coherence, illustrated by PPKED’s (Liu et al., 2021b) knowledge merging for consistency, and CA’s contrasting of abnormal features (Ma et al., 2021); (3) Cross-Modal Alignment, with (Chen et al., 2022b; Qin & Song, 2022) using attention mechanisms for text-visual feature correlation. Recent advancements (Gu et al., 2024; Tanida et al., 2023; Li et al., 2023c) focus on local visual feature learning and multi-level cross-modal alignment, showcasing the evolution of semantic matching in radiology report generation. However, these approaches mainly focus on radiology report generation, directly applying these approaches to ECG data is challenging because the patterns between medical images and ECG biomedical signals are significantly different.

### 13 Mar 2024 v2

@@ 1 Introduction
Diagnosing heart diseases is of paramount importance, requiring the analysis of multimodal data sources, which includes electrocardiogram (ECG) recordings and clinical reports. Cardiologists read and interpret these ECG recordings to manually generate comprehensive ECG reports for heart disease diagnosis, which is a complex and time-consuming process. Recently, artificial intelligence (AI) models have been developed to facilitate ECG data analysis by categorizing ECG data (Chen et al., 2022a; Hu et al., 2023; Liu et al., 2023a; Li et al., 2023b). Despite these early efforts, the automatic generation of reports from ECG recordings remains an underexplored area. Unlike other AI-empowered medical report generation applications (e.g., radiology report), the primary challenges for ECG report generation stem from the distinct nature of ECG content. ECG reports, often comprising brief phrases that summarize signal patterns, contrast with detailed anatomical descriptions in radiology reports. This difference in content and semantic interpretation between imaging and signal data complicates the direct application of radiology-focused methods to ECG reports. Furthermore, there is still a lack of comprehensive benchmarks for evaluating the performance of ECG report generation.
To tackle these challenges, we introduce MEIT, a Multimodal ECG Instruction Tuning framework that extends the capabilities of LLMs in the cardiology context to generate ECG reports using ECG signals based on human instructions. Inspired by the versatility of LLMs (Achiam et al., 2023; Touvron et al., 2023b; Wan et al., 2023b; Wang et al., 2024) in handling diverse language tasks simultaneously, we develop a specialized instruction tuning process for ECG report generation. MEIT aligns human instructions with ECG signals, enabling LLMs to generate professional-grade reports and exhibit zero-shot report generation capabilities under domain transfer scenarios across various continents and data collection devices. Specifically, leveraging publicly available ECG datasets, we construct a multimodal instruction dataset including ECG records, human instructions, and paired reports. Then, we propose an effective and efficient attention-based fusion method to integrate ECG and text representations in the latent space, enabling LLMs to understand signals for report generation without introducing additional training parameters in the attention layer.
@@ 3.2 ECG Report Generation Model
where $i$ is the current layer index. Different from gated-attention fusion in Flamingo (Alayrac et al., 2022), Memorizing Transformer (Wu et al., 2022) ,and G-MAP (Wan et al., 2022), or Q-former in BLIP-2 (Li et al., 2023a) that require the additional trainable parameters or design complex modules, we propose a simple yet effective concatenated-fusion method that directly injects the ECG embeddings with language context in the causal-attention (C-Attn) of the models. In our approach, each attention layer combines $\mathbf{H}_{e}$, generated from the ECG encoder and projector as a prefix condition, with $\mathbf{H}_{t}^{i}$, derived from the preceding layer. The fusion process is as follows:
@@ 6 Related Work
Medical Report Generation. Different from medical vision-language alignment (Wan et al., 2023a), medical Report Generation can be categorized into three main approaches: (1) Template Selection and Generation, where methods like HRGR (Li et al., 2018) focus on sentence creation or template retrieval, and CMAS (Jing et al., 2017) uses dual writers for differentiating regions; (2) Data Integration and Coherence, illustrated by PPKED’s (Liu et al., 2021b) knowledge merging for consistency, and CA’s contrasting of abnormal features (Ma et al., 2021); (3) Cross-Modal Alignment, with (Chen et al., 2022b; Qin & Song, 2022) using attention mechanisms for text-visual feature correlation. Recent advancements (Gu et al., 2024; Tanida et al., 2023; Li et al., 2023c) focus on local visual feature learning and multi-level cross-modal alignment, showcasing the evolution of semantic matching in radiology report generation. However, these approaches mainly focus on radiology report generation, directly applying these approaches to ECG data is challenging because the patterns between medical images and ECG biomedical signals are significantly different.